1.A certain gas expands from 5.768L to 9.876L at constant temperature.Calculate the work done in Joules by the gas if it expands against a constant pressure of 4.3 atm.

2.what volume of 0.050M H3PO4 solution is needed to neutralize 50mL of 0.350M KOH

3. Hydrogen gas expands from a volume of 7.65L to 10.36L under a constant pressure of 2.31atm while releasing 215Joules of heat.What is the change in energy of the system?
1
Expert's answer
2017-03-26T04:53:06-0400
1. A certain gas expands from 5.768L to 9.876L at constant temperature.Calculate the work done in Joules by the gas if it expands against a constant pressure of 4.3 atm.
Solution:
Working gas expanding
W = PdV
W= 405300 Pa x (0,009876 m^3 - 0,005768 m^3) = 1664.9724 J
Answer: 1664.9724 J

2. What volume of 0.050M H3PO4 solution is needed to neutralize 50mL of 0.350M KOH
Solution:
C1V1 = C2V2
V1 = C2V2/ C1
V1 = 50mL x 0.350M / 0.050M = 350 ml H3PO4
Answer: 350 ml H3PO4

3. Hydrogen gas expands from a volume of 7.65L to 10.36L under a constant pressure of 2.31atm while releasing 215Joules of heat.What is the change in energy of the system?
Solution:
Working gas expanding
W = PdV
We use of the first principle of thermodynamics
Q = U + W
U = Q - W
U = Q - PdV
U = 215 J - 234060,8 Pa x (0,01036 m^3- 0,00765 m^3) = - 419 J
The change in energy of the system is negative, this means that it is emitted into the environment
Answer: - 419 J